# Evaluating Large Language Models on Time Series Feature Understanding: A Comprehensive Taxonomy and Benchmark

### v1

@@ Abstract
Large Language Models (LLMs) offer the potential for automatic time series analysis and reporting, which is a critical task across many domains, spanning healthcare, finance, climate, energy, and many more. In this paper, we propose a framework for rigorously evaluating the capabilities of LLMs on time series understanding, encompassing both univariate and multivariate forms. We introduce a comprehensive taxonomy of time series features, a critical framework that delineates various characteristics inherent in time series data. Leveraging this taxonomy, we have systematically designed and synthesized a diverse dataset of time series, embodying the different outlined features. This dataset acts as a solid foundation for assessing the proficiency of LLMs in comprehending time series. Our experiments shed light on the strengths and limitations of state-of-the-art LLMs in time series understanding, revealing which features these models readily comprehend effectively and where they falter. In addition, we uncover the sensitivity of LLMs to factors including the formatting of the data, the position of points queried within a series and the overall time series length.
@@
Elizabeth Fons J.P. Morgan AI Research                        Rachneet Kaur J.P. Morgan AI Research                        Soham Palande J.P. Morgan AI Research
Zhen Zeng J.P. Morgan AI Research                        Svitlana Vyetrenko J.P. Morgan AI Research                        Tucker Balch J.P. Morgan AI Research
@@ 1 Introduction
Time series analysis and reporting play a crucial role in many areas like healthcare, finance, climate, etc. With the recent advances in Large Language Models (LLMs), integrating them in time series analysis and reporting processes presents a huge potential for automation. Recent works have adapted general-purpose LLMs for time series understanding in various specific domains, such as seizure localization in EEG time series Chen et al. (2024), cardiovascular disease diagnosis in ECG time series Qiu et al. (2023), weather and climate data understanding Chen et al. (2023), and explainable financial time series forecasting Yu et al. (2023).
@@
Despite these advancements in domain-specific LLMs for time series understanding, it is crucial to conduct a systematic evaluation of general-purpose LLMs’ inherent capabilities in generic time series understanding, without domain-specific fine-tuning. This paper aims to uncover the pre-existing strengths and weaknesses in general-purpose LLMs regarding time series understanding, such that practitioners can be well informed of areas where the general-purpose LLMs are readily applicable, and focus on areas for improvements with targeted efforts during fine-tuning.
To systematically evaluate the performance of general-purpose LLMs on generic time series understanding, we propose a taxonomy of time series features for both univariate and multivariate time series. This taxonomy provides a structured categorization of core characteristics of time series across domains. Building upon this taxonomy, we have synthesized a diverse dataset of time series covering different features in the taxonomy. This dataset is pivotal to our evaluation framework, as it provides a robust basis for assessing LLMs’ ability to interpret and analyze time series data accurately. Specifically, we examine the state-of-the-art LLMs’ performance across a range of tasks on our dataset, including time series features detection and classification, data retrieval as well as arithmetic reasoning.
Our contributions are three-fold:
@@
  • Taxonomy - we introduce a taxonomy that provides a systematic categorization of important time series features, an essential tool for standardizing the evaluation of LLMs in time series understanding.
@@
  • Diverse Time Series Dataset - we synthesize a comprehensive time series dataset, ensuring a broad representation of the various types of time series, encompassing the spectrum of features identified in our taxonomy.
  • Evaluations of LLMs - our evaluations provide insights into what LLMs do well when it comes to understanding time series and where they struggle, including how they deal with the format of the data, where the query data points are located in the series and how long the time series is.
@@ 2 Related Work
2.1 Large Language Models
Large Language Models (LLMs) are characterized as pre-trained, Transformer-based models endowed with an immense number of parameters, spanning from tens to hundreds of billions, and crafted through the extensive training on vast text datasets Zhang et al. (2024); Zhao et al. (2023). Notable examples of LLMs include Llama2 Touvron et al. (2023), PaLM Chowdhery et al. (2023), GPT-3 Brown et al. (2020), GPT4 Achiam et al. (2023), and Vicuna-13B Chiang et al. (2023). These models have surpassed expectations in numerous language-related tasks and extended their utility to areas beyond traditional natural language processing. For instance, Wang et al. (2024) have leveraged LLMs for the prediction and modeling of human mobility, Yu et al. (2023) for explainable financial time series forecasting, and Chen et al. (2024) for seizure localization. This expansive application of LLMs across diverse domains sets the stage for their potential utility in the analysis of time series data, a domain traditionally governed by statistical and machine learning models.
2.2 Language models for time series
@@
Recent progress in time series forecasting has capitalized on the versatile and comprehensive abilities of LLMs, merging their language expertise with time series data analysis. This collaboration marks a significant methodological change, underscoring the capacity of LLMs to revolutionize conventional predictive methods with their advanced information processing skills. In the realm of survey literature, comprehensive overviews provided by Zhang et al. (2024) and Jiang et al. (2024) offer valuable insights into the integration of LLMs in time series analysis, highlighting key methodologies, challenges, and future directions. Notably, Gruver et al. (2023) have set benchmarks for pre-trained LLMs such as GPT-3 and Llama2 by assessing their capabilities for zero-shot forecasting. Similarly, Xue and Salim (2023) introduced Prompcast, and it adopts a novel approach by treating forecasting as a question-answering activity, utilizing strategic prompts. Further, Yu et al. (2023) delved into the potential of LLMs for generating explainable forecasts in financial time series, tackling inherent issues like cross-sequence reasoning, integration of multi-modal data, and interpretation of results, which pose challenges in conventional methodologies. Additionally, Zhou et al. (2023) demonstrated that leveraging frozen pre-trained language models, initially trained on vast corpora, for time series analysis that could achieve comparable or even state-of-the-art performance across various principal tasks in time series analysis including imputation, classification and forecasting.
2.3 LLMs for arithmetic tasks
@@
Despite their advanced capabilities, LLMs face challenges with basic arithmetic tasks, crucial for time series analysis involving quantitative data (Azerbayev et al., 2023; Liu and Low, 2023). Research has identified challenges such as inconsistent tokenization and token frequency as major barriers (Nogueira et al., 2021; Kim et al., 2021). Innovative solutions, such as Llama2’s approach to digit tokenization (Yuan et al., 2023), highlight ongoing efforts to refine LLMs’ arithmetic abilities, enhancing their applicability in time series analysis.
@@ 3 Time Series Data
3.1 Taxonomy of Time Series Features
Our study introduces a comprehensive taxonomy for evaluating the analytical capabilities of Large Language Models (LLMs) in the context of time series data. This taxonomy categorizes the intrinsic characteristics of time series, providing a structured basis for assessing the proficiency of LLMs in identifying and extracting these features. Furthermore, we design a series of datasets following the proposed taxonomy and we outline an evaluation framework, incorporating specific metrics to quantify model performance accurately across various tasks.
Time series characteristics Description Sub-categories
@@ 3.1 Taxonomy of Time Series Features
Seasonality and Cyclical Patterns Patterns that repeat over a fixed or irregular period. Fixed-period -- constant amplitude, Fixed-period -- varying amplitude, Shifting period, Multiple seasonality
Volatility Degree of dispersion of a series over time. Constant Increasing, Clustered,
Leverage effect.
Anomalies Significant deviations from typical patterns. Spikes, step-spikes, level shifts,
@@
Statistical Properties Characteristics like fat tails, and stationarity versus non-stationarity. Fat tails, Stationarity
@@
Table 1: Taxonomy of time series characteristics.
The proposed taxonomy encompasses critical aspects of time series data that are frequently analyzed for different applications. Table 1 shows the selected features in increasing complexity, and each sub-feature. We evaluate the LLM in this taxonomy in a two-step process. In first place, we evaluate if the LLM can detect the feature, and in a second step, we evaluate if the LLM can identify the sub-category of the feature. A detailed description of the process is described in Sec. 6.1.2.
@@ 3.2 Synthetic Time Series Dataset
Leveraging our taxonomy, we construct a diverse synthetic dataset of time series, covering the features outlined in the previous section. We generated in total 9 datasets with 200 time series samples each. Within each dataset the time series length is randomly chosen between 30 and 150 to encompass a variety of both short and long time series data. In order to make the time series more realistic, we add a time index, using predominantly daily frequency. Fig. 1 showcases examples of our generated univariate time series. Each univariate dataset showcases a unique single-dimensional patterns, whereas multivariate data explore series interrelations to reveal underlying patterns. Please see Table LABEL:fig:univariate_data in the appendix for examples of each univariate dataset, and Table LABEL:fig:multivariate_data for visual examples of the multivariate cases. For a detailed description of the generation of each dataset, refer to Sec. A in the Appendix.
@@ 5.2 Performance Factors
Given a retrieval question, the position of where the queried data point occurs in the time series might impact the retrieval accuracy. Studies have discovered recency bias Zhao et al. (2021) in the task of few-shot classification, where the LLM tends to repeat the label at the end. Thus, it’s important to investigate whether LLM exhibits similar bias on positions in the task of time series understanding.
@@ 6.1.1 Models
We evaluate the following LLMs on our proposed framework: 1) GPT4. (Achiam et al., 2023) 2) GPT3.5. 3) Llama2-13B Touvron et al. (2023), and 4) Vicuna-13B Chiang et al. (2023). We selected two open-source models, Llama2 and Vicuna, each with 13 billion parameters, the version of Vicuna is 1.5 was trained by fine-tuning Llama2. Additionally we selected GPT4 and GPT3.5 where the number of parameters is unknown. In the execution of our experiments, we used an Amazon Web Services (AWS) g5.12xlarge instance, equipped with four NVIDIA A10G Tensor Core GPUs, each featuring 24 GB of GPU RAM. This setup was essential for handling both extensive datasets and the computational demands of LLMs.
@@ 6.1.2 Prompts
The design of prompts for interacting with LLMs is separated into two approaches: retrieval/arithmetic reasoning and detection/classification questioning.
@@
To evaluate the LLM reasoning over time series features, we use a two-step prompt with an adaptive approach, dynamically tailoring the interaction based on the LLM’s responses. The first step involves detection, where the model is queried to identify relevant features within the data. If the LLM successfully detects a feature, we proceed with a follow-up prompt, designed to classify the identified feature between multiple sub-categories. For this purpose, we enrich the prompts with definitions of each sub-feature (e.g. up or down trend), ensuring a clearer understanding and more accurate identification process. An example of this two-turn prompt is shown in Fig. 2. The full list can be found in Sec. F of the supplementary.
@@
We test the LLM’s comprehension of numerical data represented as text by querying it for information retrieval and numerical reasoning, as exemplified in Fig. 3 and detailed in the supplementary Sec. F.
Trend Prompts "Input:<time series>." Question 1: Detection "Question: can you detect a general upward or downward trend in this time series? Answer yes or no only." Question 2: Classification "Select one of the following answers: (a) the time series has a positive trend, (b) the time series has a negative trend. Provide your answer as either (a) or (b)."
Figure 2: Example of multi-turn prompt template used for time series feature detection and classification.
Information retrieval and arithmetic reasoning prompts "Input:<time series>. Given the input time series, please provide brief and precise answers to the following questions and format your responses in a dictionary: ’max_value’: ’Maximum value and its date.’, ’min_value’: ’Minimum value and its date.’, ’value_on_date’: ’Value of the time series on <date>’. Note: Only provide the numerical value and/or the date as the answer for each question."
Figure 3: The prompt template used for information retrieval and arithmetic reasoning evaluation.

### v2

@@ Abstract
Large Language Models (LLMs) offer the potential for automatic time series analysis and reporting, which is a critical task across many domains, spanning healthcare, finance, climate, energy, and many more. In this paper, we propose a framework for rigorously evaluating the capabilities of LLMs on time series understanding, encompassing both univariate and multivariate forms. We introduce a comprehensive taxonomy of time series features, a critical framework that delineates various characteristics inherent in time series data. Leveraging this taxonomy, we have systematically designed and synthesized a diverse dataset of time series, embodying the different outlined features, each accompanied by textual descriptions. This dataset acts as a solid foundation for assessing the proficiency of LLMs in comprehending time series. Our experiments shed light on the strengths and limitations of state-of-the-art LLMs in time series understanding, revealing which features these models readily comprehend effectively and where they falter. In addition, we uncover the sensitivity of LLMs to factors including the formatting of the data, the position of points queried within a series and the overall time series length.
@@
Elizabeth Fons  Rachneet Kaur  Soham Palande  Zhen Zeng Tucker BalchManuela VelosoSvitlana Vyetrenko {first_name}.{last_name}@jpmchase.com JP. Morgan AI Research
@@ 1 Introduction
Time series analysis and reporting are crucial in diverse fields like healthcare, finance, and climate Liu et al. (2023). The recent progress in Large Language Models (LLMs) opens exciting possibilities for automating these processes. While recent studies have explored adapting LLMs for specific time series tasks, such as seizure localization in EEG time series Chen et al. (2024), cardiovascular disease diagnosis in ECG time series Qiu et al. (2023), weather and climate data understanding Chen et al. (2023), and explainable financial time series forecasting Yu et al. (2023), a systematic evaluation of general-purpose LLMs’ inherent capabilities in understanding time series data is lacking. One notable example of domain-specific application is the BioSignal Copilot framework presented by Liu et al. (2023), which focuses on leveraging LLMs for clinical report generation from biomedical signals.
@@
This paper aims to fill this gap by uncovering the strengths and weaknesses of general-purpose LLMs in time series understanding, without any domain-specific fine-tuning. Our focus is on assessing their potential for a key downstream task: time series annotation and summarization. By understanding the baseline capabilities of LLMs, practitioners can identify areas where these models are readily applicable and areas where targeted fine-tuning efforts may be necessary to improve performance.
To systematically evaluate the performance of general-purpose LLMs on generic time series understanding, we propose a taxonomy of time series features for both univariate and multivariate time series. This taxonomy serves as a structured framework for evaluating LLM performance and provides a foundation for future research in this domain. Based on this taxonomy, we have created a diverse synthetic dataset of time series that covers a wide range of features, each accompanied by qualitative and quantitative textual descriptions.
Our evaluations focus on tasks directly relevant to time series annotation and summarization, such as feature detection, classification, and data retrieval as well as arithmetic reasoning. Additionally, we assess the LLMs’ ability to match textual descriptions to their corresponding time series, leveraging the textual descriptions in our dataset. These findings will be instrumental for developing LLM-powered tools for automated time series annotation and summarization, ultimately enhancing data analysis and reporting workflows across diverse domains. Our contributions are three-fold:
@@
  • Taxonomy - we introduce a comprehensive taxonomy that provides a systematic categorization of important time series features, an essential tool for standardizing the evaluation of LLMs in time series understanding.
@@
  • Diverse Time Series Dataset - we synthesize a diverse time series dataset with train/validation/test splits, ensuring a broad representation of various time series types, encompassing the spectrum of features identified in our taxonomy, each with accompanying textual descriptions.
  • Evaluations of LLMs - our evaluations provide insights into LLMs’ strengths and weaknesses in understanding time series. We analyze how LLMs handle data format, query location, and time series length, providing a nuanced understanding of their capabilities in this domain.
@@ 2 Related Work
Large Language Models
Large Language Models (LLMs), such as Llama2 Touvron et al. (2023), PaLM Chowdhery et al. (2023), GPT-3 Brown et al. (2020), GPT4 Achiam et al. (2023), and Vicuna-13B Chiang et al. (2023), have demonstrated remarkable capabilities in various language-related tasks and have recently been explored for their potential in time series analysis.
Language Models for Time Series
@@
Recent progress in time series forecasting has capitalized on the versatile and comprehensive abilities of LLMs, merging their language expertise with time series data analysis. This collaboration marks a significant methodological change, underscoring the capacity of LLMs to revolutionize conventional predictive methods with their advanced information processing skills. Notably, Gruver et al. (2023) have set benchmarks for pre-trained LLMs such as GPT-3 and Llama2 by assessing their capabilities for zero-shot forecasting. Similarly, Xue and Salim (2023) introduced Prompcast, adopting a novel approach by treating forecasting as a question-answering activity, utilizing strategic prompts. Further, Yu et al. (2023) delved into the potential of LLMs for generating explainable forecasts in financial time series, tackling inherent issues like cross-sequence reasoning, integration of multi-modal data, and interpretation of results, which pose challenges in conventional methodologies. Additionally, Zhou et al. (2023) demonstrated that leveraging frozen pre-trained language models, initially trained on vast corpora, for time series analysis could achieve comparable or even state-of-the-art performance across various principal tasks in time series analysis including imputation, classification and forecasting.
Recent advancements in the application of LLMs to biomedical time series data have also shown promise in the automated generation of clinical reports. Liu et al. (2023) introduce BioSignal Copilot, a system that leverages LLMs for drafting reports from biomedical signals, such as electrocardiograms (ECGs) and electroencephalograms (EEGs). Their work highlights the importance of domain-specific feature extraction in facilitating LLM understanding of time series data, aligning with our work on developing a comprehensive taxonomy of time series features to enhance LLM interpretability and analysis in various applications. Notably, their focus on automatic report generation from the processed signals serves as a specific downstream task, further emphasizing the need for a systematic evaluation of LLMs’ ability to understand and extract relevant features from time series data, such as the one presented in this work.
LLMs for arithmetic tasks
@@
Despite their advanced capabilities, LLMs face challenges with basic arithmetic tasks, crucial for time series analysis involving quantitative data (Azerbayev et al., 2023; Liu and Low, 2023). Research has identified challenges such as inconsistent tokenization and token frequency as major barriers (Nogueira et al., 2021; Kim et al., 2021). Innovative solutions, such as Llama2’s approach to digit tokenization Yuan et al. (2023), highlight ongoing efforts to refine LLMs’ arithmetic abilities, enhancing their applicability in time series analysis.
@@ 3 Time Series Data
Table 1: Taxonomy of time series characteristics.
Main Category Description Sub-category
@@
Seasonality and Cyclical Patterns Patterns that repeat over a fixed or irregular period. Fixed-period, Shifting period, Multiple seasonality
Anomalies Significant deviations from typical patterns. Spikes, level shifts,
@@
Volatility Degree of dispersion of a series over time. Constant, Trending, Clustered,Dynamic
@@
Stationarity Properties Stationarity versus non-stationarity. Stationarity
Distribution Properties Characteristics like fat tails Fat tails
@@
3.1 Taxonomy of Time Series Features
Our study introduces a comprehensive taxonomy for evaluating the analytical capabilities of Large Language Models (LLMs) in the context of time series data. This taxonomy categorizes the intrinsic characteristics of time series, providing a structured basis for assessing the proficiency of LLMs in identifying and extracting these features. The proposed taxonomy encompasses critical aspects of time series data that are frequently analyzed for different applications and are commonly used in qualitative descriptions of time series data. These features are considered the most relevant for evaluating the ability of LLMs to generate and understand textual reports of time series data.
The features are organized in increasing order of complexity, starting with trend, seasonality, volatility, anomalies, structural breaks, and distribution properties. Each main feature is further divided into sub-categories to provide a more nuanced evaluation of LLM capabilities. This hierarchical organization allows for a detailed assessment of LLM performance on both simple and complex time series characteristics. Table 1 presents the selected features in order of increasing complexity and their sub-features. While we have strived to define the features as distinctly as possible, it is important to note that some overlap may exist between certain categories.
Justification for the proposed taxonomy
Our selection of features is based on extensive literature review and expert consultations. Trends and seasonality are fundamental components widely recognized in time series analysis across various domains, such as finance and climate science (Hyndman and Athanasopoulos, 2018; Shumway and Stoffer, 2000). Volatility and anomalies are crucial for understanding dynamic behaviors and identifying significant deviations in data (Tsay, 2005; Chandola et al., 2009). Structural breaks and distribution properties are essential for capturing shifts in underlying data generation processes and understanding the statistical nature of the data (Perron, 2005; Cont, 2001). Table 5 provides definitions of each sub-category along with domain examples where these features could be referenced.
@@ 3.2 Synthetic Time Series Dataset
Leveraging our taxonomy, we construct a diverse synthetic dataset of time series, covering the features outlined in the previous section. We generated in total 10 datasets, each with a training split (5000 samples), validation split (2000 samples), and test split (200 samples) to facilitate model development and evaluation. Within each dataset, the time series length is randomly chosen between 30 and 150 to encompass a variety of both short and long time series data. In order to make the time series more realistic, we add a time index, using predominantly daily frequency. Each time series in the dataset is accompanied by a qualitative description, a textual summary of the main features present in the time series (e.g., "This time series exhibits a downward quadratic trend, commencing with higher figures and falling gradually."), and a quantitative description, which includes the minimum and maximum values, the date range, and a textual description of the specific features present (e.g., "This daily time series covers the period from 2024-01-01 to 2024-05-04. It exhibits multiple seasonal patterns with monthly seasonality, with 5 peaks and 4 troughs, and an average amplitude of 24.25."). Fig. 1 showcases examples of our generated univariate time series. Each univariate dataset showcases a unique single-dimensional pattern, whereas multivariate data explore series interrelations to reveal underlying patterns. See Table LABEL:fig:univariate_data and Table LABEL:fig:multivariate_data in the appendix for visual examples of each dataset. For a detailed description of the generation of each dataset, refer to Appendix. B.
@@ 4 Time Series Benchmark Tasks
Time Series Description
To evaluate the ability of LLMs to match time series to their corresponding descriptions, even in the presence of distractors, we introduce two new tasks: (1) Text Matching (inter-dataset): the LLM is presented with a time series and four different descriptions from the same dataset, one of which is the correct description for the given time series. The descriptions include both qualitative commentaries and quantitative information about the time series. The LLM is asked to select the description that is closest to the time series. This task assesses the LLM’s ability to match a time series to its corresponding description, even in the case where the qualitative description is similar; (2) Text Matching (cross-dataset): the LLM is presented with a time series and four different qualitative descriptions, each from a different dataset. This task assesses the LLM’s ability to match a time series to its corresponding description based only on qualitative features, without relying on any quantitative information.
Metric GPT4 GPT3.5 Llama2 Vicuna Phi3
Zero-shot CoT Zero-shot CoT Zero-shot CoT Zero-shot CoT Zero-shot CoT
Univariate time series characteristics
Feature detection
Trend F1score 0.79 0.89 0.45 0.66 0.51 0.56 0.58 0.58 0.72 0.78
Seasonality F1score 0.94 0.98 0.43 0.55 0.64 0.35 0.49 0.48 0.82 0.83
Anomalies F1score 0.84 0.81 0.57 0.47 0.47 0.51 0.49 0.52 0.43 0.71
Volatility F1score 0.68 0.73 0.43 0.43 0.42 0.53 0.45 0.48 0.73 0.69
Struct. break F1score 0.59 0.61 0.57 0.48 0.39 0.44 0.48 0.52 0.44 0.67
Stationarity F1score 0.33 0.59 0.33 0.40 0.33 0.39 0.44 0.42 0.33 0.46
Fat Tails F1score 0.39 0.44 0.36 0.34 0.39 0.44 0.48 0.47 0.45
Feature classification
Trend F1score 0.98 0.98 0.78 0.95 0.43 0.70 0.53 0.48 0.48 0.95
Seasonality F1score 0.17 0.21 0.17 0.16 0.31 0.27 0.23 0.18 0.48 0.24
Anomalies F1score 0.87 0.95 0.20 0.40 0.30 0.37 0.37 0.44 0.53 0.48
Volatility F1score 0.18 0.25 0.07 0.16 0.12 0.10 0.15 0.17 0.08 0.15
Struct. break F1score 0.42 0.41 0.56 0.57 0.30 0.43 0.41 0.35 0.51 0.47
Multivariate time series characteristics
Fixed Corr. F1score 0.48 0.39 0.43 0.38 0.43 0.40 0.46 0.43 0.57
Lagged Corr. F1score 0.54 0.52 0.46 0.45 0.42 0.42 0.45 0.41 0.40
Changing Corr. F1score 0.48 0.43 0.44 0.52 0.43 0.50 0.45 0.48 0.65
Information Retrieval
Value on Date Acc 1.00 1.00 0.99 0.99 0.54 0.49 0.61 0.62 0.93 0.89
Value on Date MAPE 0.00 0.00 0.03 0.03 1.06 0.73 0.75 0.76 0.19 0.17
Arithmetic Reasoning
Min Value Acc 1.00 0.99 0.99 0.98 0.63 0.55 0.63 0.72 0.94 0.91
MAPE 0.00 0.00 0.01 0.01 3.89 7.42 3.96 4.70 0.10 0.41
Min Date Acc 0.98 0.94 0.93 0.93 0.40 0.32 0.42 0.49 0.85 0.82
Max Value Acc 1.00 1.00 0.96 0.94 0.53 0.54 0.47 0.57 0.87 0.78
MAPE 0.00 0.00 3.66 3.96 3.23 1.09 3.12 2.27 0.11 0.26
Max Date Acc 0.99 0.93 0.91 0.90 0.32 0.34 0.29 0.37 0.77 0.70
@@ 5.2 Performance Factors
Given a retrieval question, the position of where the queried data point occurs in the time series might impact the retrieval accuracy. Studies have discovered recency bias Zhao et al. (2021) in the task of few-shot classification, where the LLM tends to repeat the label at the end. Thus, it is important to investigate whether LLM exhibits similar bias on positions in the task of time series understanding.
@@ 6.1.1 Models
We evaluate the following LLMs on our proposed framework using the test split of our dataset: 1) GPT4. (Achiam et al., 2023) 2) GPT3.5. 3) Llama2-13B Touvron et al. (2023), 4) Vicuna-13B Chiang et al. (2023), and 5) Phi3-Medium (14B)et al. (2024). We selected three open-source models, Phi3, Llama2 and Vicuna, the first with 14B parameters and the remaining with 13 billion; the version of Vicuna is 1.5 and was trained by fine-tuning Llama2. Additionally we selected GPT4 and GPT3.5 where the number of parameters is unknown. In the execution of our experiments, we used an Amazon Web Services (AWS) g5.12xlarge instance, equipped with four NVIDIA A10G Tensor Core GPUs, each featuring 24 GB of GPU RAM.
@@ 6.1.2 Prompts
The design of prompts for interacting with LLMs is separated into two approaches: retrieval/arithmetic reasoning and detection/classification questioning. In addition to zero-shot prompting, we also use chain-of-thought (CoT) Wei et al. (2022) prompting to enhance the reasoning capabilities of LLMs. We employ regular expressions to parse the responses for feature detection and classification tasks in the zero-shot setting. However, for chain-of-thought prompting, we utilize an LLM to parse the responses due to their increased complexity and length.
@@
To evaluate the LLM reasoning over time series features, we use a two-step prompt with an adaptive approach, dynamically tailoring the interaction based on the LLM’s responses. The first step involves detection, where the model is queried to identify relevant features within the data. If the LLM successfully detects a feature, we proceed with a follow-up prompt, designed to classify the identified feature between multiple sub-categories. For this purpose, we enrich the prompts with definitions of each sub-feature (e.g. up or down trend), ensuring a clearer understanding and more accurate identification process. The full list of prompts can be found in Sec. G of the supplementary.
@@
We test the LLM’s comprehension of numerical data represented as text by querying it for information retrieval and numerical reasoning, as exemplified in Fig. LABEL:fig:ret_prompt and detailed in the supplementary Sec. G.